Rhesus factor, is a type of protein found on the outside of red blood cells. The protein is genetically inherited. A donor is said to have Rh+ blood if the Rhesus factor is present in their blood. Suppose that a blood donation campaign is launched by the University of Ottawa walk-in clinic. Among undergraduate students donating blood to walk-in clinic, 30% have Rh+.

a) If seven undergraduate students donate blood at this clinic on a particular day, what is the probability that at least three of them do not have the Rh factor? What is the probability that the number of students with Rh+ blood will be within 0.1 of the true population mean?

b) How many undergraduate students must donate blood to have the probability of at least one Rh+ donor over 0.90?
Answer #1

a)

P(X>=3)

= 1 - P(X<=2)

X P(X)
P ( X = 0) = C (7,0) * 0.3^0 * ( 1 - 0.3)^7= 0 0.0824
P ( X = 1) = C (7,1) * 0.3^1 * ( 1 - 0.3)^6= 1 0.2471
P ( X = 2) = C (7,2) * 0.3^2 * ( 1 - 0.3)^5= 2 0.3177

= 0.3529

b)

For 7 students

X P(X)
P ( X = 0) = C (7,0) * 0.3^0 * ( 1 - 0.3)^7= 0 0.0824

= 1 - P(X=0)

= 0.9176

Please let me know in case of any doubt.

Thanks in advance!


Please upvote!